HETAV PATEL

Grade 11,

Convert each part of the ratio to percentage: 1:4 Convert each part of the ratio to percentage: 1:4

Convert each part of the ratio to percentage: 1:4 
Convert each part of the ratio to percentage: 1:4 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to find total parts by adding the given ratio = 1 + 4 = 5
1stpart = (1/5) = (1/5) × 100 %
= 1 × 20%
= 20%
2ndpart = (4/5)
= (4/5) × 100%
= 4 × 20
= 80%

Thanks